OpenStudy (anonymous):

HELPPP!!! Find a point c satisfying the conclusion of the MVT for the following function and interval (Round your answer to three decimal places.) f(x) = x^−3, [1, 3] i've gotten that (-8/9)/2

OpenStudy (amistre64):

whats the slope from 1 to 3?

OpenStudy (anonymous):

umm... It doesn't give that I don't think

OpenStudy (amistre64):

this for integration or derivatives

OpenStudy (anonymous):

oh is it -3/c^4

OpenStudy (amistre64):

depends; what is the application of the MVT in this case?

OpenStudy (amistre64):



OpenStudy (anonymous):

that is it, because the example problem has that to the left, its like this -2/c^3 = f ' (c) = f(b)- f(a) / b-a

OpenStudy (amistre64):

ok, then its the derivative one

OpenStudy (anonymous):

so this one would be -3/c^4 = f ' (c) = f(b)-f(a)/ b-a

OpenStudy (amistre64):

whats the slope of the line from 1 to 3?

OpenStudy (anonymous):

idk im guessing it is -3/c^4 the derivate of x^-3 with f '(c)

OpenStudy (amistre64):

\[slope =\frac{f(3)-f(1)}{3-1}\]

OpenStudy (amistre64):

the change in y over the change in x

OpenStudy (anonymous):

oh it is (-8/9) / 2

OpenStudy (amistre64):

\[\cfrac{\frac{1}{27}-\frac{1}{1}}{2}=\cfrac{\frac{1-27}{27}}{2}=\frac{-26/27}{2/1}=-13/27\]

OpenStudy (amistre64):

so, at what point "c" is the derivative equal to -13/27

OpenStudy (anonymous):

ahh dang, i see, and then you somehow plug that into -3/c^4 im guessing?

OpenStudy (amistre64):

yes, you then equate that slope with the derivative at "c"

OpenStudy (amistre64):

\[\frac{-3}{c^4}=\frac{-13}{27}\]

OpenStudy (anonymous):

-13/-9 = c^4

OpenStudy (anonymous):

wait -9/13 = c^4

OpenStudy (amistre64):

\[\frac{-3(27)}{-13}=c^4\to \color{#ff0080} {c=\sqrt[4]{\frac{-3(27)}{-13}}}\]

OpenStudy (anonymous):

holy crap haha, now the question is what is that in decimal form

OpenStudy (amistre64):

whatever the calculator gives you :)

OpenStudy (amistre64):

c=(-3(27/(-13)))^(1/4)

OpenStudy (anonymous):

my calculator gae me 1.580

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

thanks!!

OpenStudy (amistre64):

yw

OpenStudy (anonymous):

1/2 c ^-1/2 = 1/9 is that the same as 1/c^2 = 1/9

OpenStudy (amistre64):

\[\frac{\sqrt{c}}{2}\ne c^{-2} \text{ in general}\]

OpenStudy (anonymous):

y = \[\sqrt{x}\]

OpenStudy (anonymous):

so c would equal something weird

OpenStudy (amistre64):

i dont know, depends on what you are trying to do ...

OpenStudy (anonymous):

the derivative of the square root of x = 1/9

OpenStudy (amistre64):

(4x)^{-1/2} = 9^{-1} hmmm

OpenStudy (amistre64):

((4x)^{-1/2} = 9^{-1} )^{-2} 4x = 9^2 x = 81/4

OpenStudy (amistre64):

maybe

OpenStudy (amistre64):

yeah, thats good lol

OpenStudy (anonymous):

so you did 1/2 (x) ^-1/2 = 2*2x ^-1/2 = 1/9 so 9^-1 = 4x^-1/2 so then you square 9 and divide by 4

OpenStudy (amistre64):

i did yes; but its simpler than that really :) \[2\sqrt{x}=9\] \[\sqrt{x}=9/2\] \[x=81/4\]

OpenStudy (anonymous):

oh okay gotcha (: i need help with another one, it involves sin and cosine

OpenStudy (amistre64):

<---- post it to the left

OpenStudy (anonymous):

okayyy